OpenStudy (anonymous):

The total number of people infected with a virus often grows like a logistic curve. Suppose that 10 people originally have the virus, and that in the early stages of the virus (with time, t , measured in weeks), the number of people infected is increasing exponentially with K=1.6 . It is estimated that, in the long run, approximately 6750 people become infected. Use this information to find a logistic function to model this situation. P=? The total number of people infected with a virus often grows like a logistic curve. Suppose that 10 people originally have the virus, and that in the early stages of the virus (with time, t , measured in weeks), the number of people infected is increasing exponentially with K=1.6 . It is estimated that, in the long run, approximately 6750 people become infected. Use this information to find a logistic function to model this situation. P=? @Mathematics

OpenStudy (anonymous):

6750 because if it has a long time to spread it eventually gets everyone

OpenStudy (anonymous):

so do i do 10=1650/1+be^(k(0)) then solve for b?

OpenStudy (anonymous):

you have to use the P=L/(1+be^(-kt))

OpenStudy (anonymous):

somehow

OpenStudy (anonymous):

it doesn't specify the time in weeks when this occurs

OpenStudy (valpey):

\[\frac{dP}{dt}=1.6P\left(1 - \frac{P}{6750}\right)\]

OpenStudy (anonymous):

right it is a differential equation

OpenStudy (anonymous):

Valpey can you continue?

OpenStudy (anonymous):

sure we have to separate the variables all the p's on one side and the dt term on the right. We then integrate both sides.

OpenStudy (anonymous):

separable first order

OpenStudy (anonymous):

i got (6750)/1+674e^-(1.6t)

OpenStudy (anonymous):

i got it right